OpenStudy (anonymous):

20,000 people voted in each county, and there are 7 counties. What is the order of magnitude for the total number of people who voted?

OpenStudy (anonymous):

7*20,000 = 140,000. Order of magnitude is number of digits minus 1. 6 - 1 = 5. There order of magnitude is 5.

OpenStudy (anonymous):

1.4 * 10^5 = 140,000 is another way to look at it.